### v1

@@ 5 Implementation
We developed a command-line Python application for our approach, allowing users to interactively configure the process, manipulate the data, and assess the mining algorithm’s output. The scripts we created are available at https://github.com/unibz-core/CM-Mining.
@@
For tasks involving graph importing, processing, and transformation, we used the NetworkX library444https://networkx.org/, a powerful Python package for creating, analyzing, and visualizing complex networks. NetworkX played a crucial role in each step of the entire workflow. We integrated the Grandiso library555https://pypi.org/project/grandiso/—a versatile Python library known for its efficient graph-matching and isomorphism-checking capabilities—in tasks involving graph matching, isomorphism detection, and motif search. For pattern mining, we relied on the gSpan Python implementation666https://pypi.org/project/gspan-mining/ of the frequent subgraph mining algorithm. This library enabled us to discover recurring subgraph patterns within our data. We achieved the visualization of discovered patterns using PlantUML777https://plantuml.com/, a flexible tool for generating UML diagrams. PlantUML seamlessly integrated into our framework, enhancing the interpretability of our results through clear and informative visual representations.
This suite of tools enabled a robust and efficient framework for structural pattern discovery and visualization. These tools collectively ensured the accuracy, scalability, and comprehensibility of our approach, enabling us to extract valuable knowledge from complex conceptual models.
5.1 Ontology-Driven Conceptual Modeling
@@
To establish our scientific contribution’s proof of concept, we focused our implementation, in particular by adapting our importing step, on UML class diagrams, specifically on OntoUML models. These represent the output of a trending paradigm situated at the confluence of conceptual modeling and ontology engineering, namely Ontology-driven conceptual modeling (ODCM). ODCM frequently entails the utilization of fundamental ontologies to steer the formulation of conceptual models, modeling languages, and tools [34]. Within this framework, the OntoUML modeling language [35, 36], has risen to prominence as one of the predominant methodologies [34].
@@
OntoUML extends the UML modeling language. Its meta-model is grounded on UFO (Unified Foundational Ontology) [37], a formal theory based on contributions from Formal Ontology in Philosophy, Philosophical Logic, Cognitive Psychology, and Linguistics. UFO is one of the most used foundational ontologies in conceptual modeling and OntoUML is among the most used languages in ontology-driven conceptual modeling [34]. An example of core ontological distinctions underlying OntoUML is represented by the key categories of object types (e.g., Kind, Subkind, Roles and RoleMixins), trope types (e.g., Relator, Mode) and Relations (FormalRelations, MaterialRelations and ParthoodRelations).888For an in-depth analysis and characterization of the ontological categories underlying OntoUML, the reader is referred to [35, 36].
@@ 6 Experiments
Can the proposed approach identify previously recognized interesting patterns? This research question is aimed at testing whether the proposed solution can discover pre-identified interesting patterns (R1). This research question is also used to check the role of the customization steps in supporting the discovery process (R2) and the level of comprehensibility of the outcome (R3).
@@
Is the clustering step accurate in grouping patterns? This research question is mainly concerned with (R3). Here, we want to assess the practical utility of a key component in the presentation of the output.
The data about the experiments and detailed instructions for reproducibility are available at this GitHub link: unibz-core/CM-Mining_ExperimentData.
@@ 6.1 Experiment 1: Reliability Test
Data: A dataset comprising 10 models developed by our team. These models were created specifically for this experiment, taking inspiration from structures that are present in real models, with the goal of reproducing a controlled number of target patterns. The distribution of patterns per model was as from Table 1.
Setup: For validation, we used six common OntoUML patterns that served as “litmus test”.999A litmus test is “a critical indicator of future success or failure” A is a litmus test for B if A can be effectively used to measure some property of B. [38] These patterns were previously manually identified as useful for building OntoUML models by the designers of the language within multiple example models [7, 39]. We represent the selected patterns in Figure 4.
We executed the application seven times (one without adopting parameters to filter specific concepts and the other times with six different configurations, each to find one of the 6 selected target patterns) and we checked whether the proposed solution could discover the pre-identified interesting patterns (R1). Moreover, we tested the role of the customization steps in supporting the discovery process (R2) and the level of comprehensibility of the outcome (R3). First, we conducted a trial adopting no customization facility from the pipeline. As parameters, we selected 3 (the support of the less frequent patterns, i.e., Relator, Subkind and Phase) as minimum support, and 5 as minimum number of nodes (the size of the smaller pattern, i.e., Characterization),101010Where we have 2 nodes for the classes, 1 node as an association and 2 nodes for the cardinalities. so that all patterns can be discovered. For each other trial, we customized the pipeline to find the target pattern, e.g., by filtering some constructs and some other patterns we did not want to discover. For instance, for the Relator pattern, we selected Kind, Relator, and Role as target constructs and we filtered out information about the generalizations sets (the customization we adopted in terms of minimum support, nodes number and selected/removed constructs is reported in Table 2.).
@@
trial | freq. | nodes | filter | patterns |
@@
Relator | 3 | 12 | Select: ‘kind’, ‘role’, ‘relator’. Remove: ‘charact’ | 42 |
RoleMixin | 4 | 10 | Select: ‘kind’, ‘role’, ‘rmixin’, ‘genset’. Remove: ‘assoc’ | 2 |
Charact. | 8 | 4 | Select: ‘kind’, ‘mode’. Remove: ‘gen’ | 1 |
Category | 6 | 4 | Select: ‘kind’, ‘category’, ‘genset’. Remove: ‘assoc’ | 32 |
Subkind | 3 | 4 | Select: ‘kind’, ‘subkind’, ‘gen-set’. Remove: ‘assoc’ | 41 |
Phase | 3 | 4 | Select: ‘kind’, ‘phase’, ‘genset’. Remove: ‘assoc’ | 89 |
Results: The conducted trial identified the target patterns with 100% accuracy, demonstrating the pipeline’s proficiency in addressing R1. The approach identified all the (classes of) patterns that were explicitly identified a priori and correctly counted their occurrences.
@@
Threats to validity: This experiment focused on a limited set of patterns that, although relatively intricate, represent only a fraction of a broader spectrum of patterns. Moreover, the synthetic dataset used in this study was fairly compact. However, the combination of the patterns we selected is suitable to assess whether our approach can discover structures with the level of expressiveness we require.
@@ 6.2 Experiment 2: Performance Test
Setup: This validation was executed on a MacBook Pro (Retina, 13-inch, Early 2015) with CPU 2,7 GHz Intel Core i5 and was organized in two main trials: a baseline trial (0), where we selected as input the 47 conceptual models from the OntoUML dataset and trial (1) where we added 47 more models. The selected models have different sizes, in terms of relations and and classes. As an example, the graphs generated from these sources can have a small number of nodes and relations, namely 47 and 35 respectively (see lindeberg2022simple-ontorights.json, in the reference experiments repository), or higher, namely 2449 and 2282 (see indeberg2022full-ontorights.json). For each step, to increase the amount of data to be handled, we also tested the approach with six different customizations (we show these values in Table 3) concerning partition size parameters and the minimum frequency of the mining task.
@@
models | filtering (s) | nodes | freq. | mining (s) | patterns | clustering (s) | viz. (s) |
47 | 0,7327 | 5 | 20 | 1,375 | 1 | 0,011 | 0,106 |
3 | 20 | 1,591 | 15 | 0,043 | 1,648 | ||
5 | 15 | 3,792 | 22 | 0,073 | 2,257 | ||
3 | 15 | 3,950 | 65 | 0,422 | 6,669 | ||
5 | 10 | 47,579 | 246 | 0,982 | 16,387 | ||
3 | 10 | 48,448 | 372 | 1,360 | 20,387 | ||
94 | 1,238 | 5 | 20 | 130,458 | 37 | 0,103 | 3,781 |
3 | 20 | 184,239 | 72 | 0,503 | 7,360 | ||
5 | 15 | 224,782 | 141 | 0,734 | 9,410 | ||
3 | 15 | 229,138 | 193 | 0,812 | 11,500 | ||
5 | 10 | 324,781 | 586 | 1,990 | 25,675 | ||
3 | 10 | 359,318 | 662 | 3,156 | 31,057 |
Threats to validity: The primary challenge to the validity of this experiment stems from the absence of a comparative analysis involving different devices. The ultimate performance might be influenced by additional variables not present in the current setting. Nonetheless, the configuration we employed can be viewed as a stress test, given its resemblance to the common features found in widely utilized laptops and the absence of a high-performance CPU.

### v2

@@ 5 Implementation
We developed a command-line Python application for our approach, allowing users to interactively configure the process, select the input conceptual modeling language, manipulate the data, and assess the mining algorithm’s output. The scripts we created are available at https://github.com/unibz-core/CM-Mining.
@@
For tasks involving graph importing, processing, and transformation, we used the NetworkX library444https://networkx.org/, a powerful Python package for creating, analyzing, and visualizing complex networks. NetworkX played a crucial role in each step of the entire workflow. We integrated the Grandiso library555https://pypi.org/project/grandiso/—a versatile Python library known for its efficient graph-matching and isomorphism-checking capabilities—in tasks involving graph matching, isomorphism detection, and motif search. For pattern mining, we relied on the gSpan Python implementation666https://pypi.org/project/gspan-mining/ of the frequent subgraph mining algorithm. This library enabled us to discover recurring subgraph patterns within our data. We achieved the visualization of discovered patterns using PlantUML777https://plantuml.com/, a flexible tool for generating UML and ArchiMate diagrams. PlantUML is seamlessly integrated into our framework, enhancing the interpretability of our results through clear and informative visual representations.
This suite of tools enabled a robust and efficient framework for graph structure discovery and visualization, ensuring we can efficiently access useful information from the mining outputs.
5.1 OntoUML Miner
@@
As an initial proof of concept for our scientific contribution, we adapted our pipeline to support UML class diagrams, focusing specifically on OntoUML models. These represent the output of a trending paradigm situated at the confluence of conceptual modeling and ontology engineering, namely Ontology-driven conceptual modeling (ODCM). ODCM frequently entails the utilization of fundamental ontologies to steer the formulation of conceptual models, modeling languages, and tools [38]. Within this framework, the OntoUML modeling language [39, 40], has risen to prominence as one of the predominant methodologies [38].
Data: Set of of Conceptual Models Result: Set of Labeled Property Graphs (LPG) 1 for every conceptual model , where do 2 map each class into a node ; 3 map each association into a a node ; 4 map each generalization into a node ; 5 map association cardinalities into a node ; 6 for each generalization node do 7 Connect to the parent class-node with an edge “general”; 8 Connect to the child class-node with an edge “specific”; 9 end for 10 for each association node do 11 Connect to the source class-node with an edge “source”; 12 Connect to the target class-node with an edge “target”; 13 for each cardinalities node do 14 Connect to the corresponding with an edge “cardinalities”; 15 end for 16 17 end for 18 for each class do 19 create the corresponding label(s) for node ; 20 end for 21 for each association do 22 create the corresponding label(s) for node ; 23 24 end for 25 create ; 26 add a graph to ; 27 28 end for 29
@@
OntoUML extends the UML modeling language. Its meta-model is grounded on UFO (Unified Foundational Ontology) [41], a formal theory based on contributions from Formal Ontology in Philosophy, Philosophical Logic, Cognitive Psychology, and Linguistics. UFO is one of the most used foundational ontologies in conceptual modeling and OntoUML is among the most used languages in ontology-driven conceptual modeling [38]. An example of core ontological distinctions underlying OntoUML is represented by the key categories of object types (e.g., Kind, Subkind, Roles, and RoleMixins), trope types (e.g., Relator, Mode) and Relations (FormalRelations, MaterialRelations, and ParthoodRelations).888For an in-depth analysis and characterization of the ontological categories underlying OntoUML, the reader is referred to [39, 40].
@@
The purpose of the OntoUML miner is to enable the discovery of recurrent structures within OntoUML models, requiring the ability to handle all OntoUML constructs. To achieve this, the module adapts both the importing and visualization steps of the pipeline, which were identified as language-dependent in Figure 4. The OntoUML-specific importing step is detailed in Algorithm 3, providing a specialized version of Algorithm 1. This partial view of the algorithm shows that the input concepts from the conceptual model correspond to specific OntoUML constructs, many of which also appear in UML. For example, taxonomic relations, associations, and relation properties like cardinalities are represented.999Note that the example pseudocode presented covers only a subset of constructs. Information on generalization sets, aggregation, and composition relations is not included. However, the importing step can readily handle these elements by applying the same reification strategy outlined in Algorithm 1 (e.g., a generalization set is represented as a node linked to generalization relation nodes, with properties such as disjoint and complete).
As a final remark on the OntoUML visualization step, this process entails adapting the PlantUML transformation so that output patterns are displayed using OntoUML-like notation.
5.2 ArchiMate Miner
Besides OntoUML, we extended our pipeline to support ArchiMate models, demonstrating the adaptability of our mining approach to different modeling languages. ArchiMate, standardized by the Open Group,101010https://pubs.opengroup.org/architecture/archimate3-doc/ is one of the most widely used Enterprise Architecture (EA) modeling language [42]. The ArchiMate framework adopts a layered view of an enterprise, where the core entities of an enterprise are categorized along layers (e.g., Business, Application, or Technology) and aspects (e.g., Active Structure, Passive Structure, or Behavior).
To enable the discovery of frequent structures within ArchiMate models, we adapted our pipeline to accommodate ArchiMate constructs during the language-dependent importing, filtering, and visualization steps.
The ArchiMate-specific importing step is detailed in Algorithm 4, extending the general approach outlined in Algorithm 1. This step transforms ArchiMate models into Labeled Property Graphs (LPGs) by mapping elements and relationships to nodes and assigning labels that capture their name and type information (e.g., BusinessProcess for an element or Assignment for a relationship). Specialization relationships are treated distinctly, with nodes connected by edges labeled general and specific to ensure that hierarchical structures are distinctly represented and allow for additional filtering.
The filtering step allows users to refine the models before mining, reducing the graph size and, subsequently, the search space. Users can apply filters based on:
.i
element types (e.g., ApplicationComponent, TechnologyService),
.ii
layers or aspects (e.g., focus solely on the Technology Layer or Behavior Aspect),
.iii
relationship types (e.g., include only Realization or Access relationships), and
.iv
edge labels (e.g., filtering by general/specific or source/target labels).
Finally, similar to the OntoUML implementation, the visualization step was adapted to render ArchiMate patterns using PlantUML, which natively supports ArchiMate diagrams.111111https://plantuml.com/archimate-diagram
Data: Set of of ArchiMate Models Result: Set of Labeled Property Graphs (LPG) 1 for every model , where do 2 map each element into a node ; 3 map each relationship into a a node ; 4 map each specialization into a node ; 5 for each specialization node do 6 Connect to the parent element-node with an edge “general”; 7 Connect to the child element-node with an edge “specific”; 8 end for 9 for each relationship node do 10 Connect to the source element-node with an edge “source”; 11 Connect to the target element-node with an edge “target”; 12 13 end for 14 for each element do 15 create the corresponding label(s) for node ; 16 end for 17 for each relationship do 18 create the corresponding label(s) for node ; 19 20 end for 21 create ; 22 add graph to ; 23 24 end for
@@ 6 Experiments
Can the proposed approach generate structures encoding previously recognized interesting patterns? This research question is aimed at testing whether the proposed solution can discover pre-identified interesting patterns (R1). This research question is also used to check the role of the customization steps in supporting the discovery process (R2) and the level of comprehensibility of the outcome (R3).
@@
Is the clustering step accurate in grouping structures? This research question is mainly concerned with (R3). Here, we want to assess the practical utility of a key component in the presentation of the output.
The data about the experiments and detailed instructions for reproducibility are available at this GitHub link: https://github.com/unibz-core/cmining-approach.
@@ 6.1 Experiment 1: Reliability Test
The primary objective of this experiment is to validate the approach with respect to RQ1. Specifically, we aim to determine whether the choices made in the importing task enable the discovery of structures similar to those previously identified by domain experts. To achieve this, we used known patterns for each language as a reference and observed the impact of selected input parameters on the mining task outputs. For full control over the validation, we created ten models, each containing instances of the known patterns. This controlled context and small dataset also allowed us to closely monitor how filtering actions effectively prune results that fall outside the search scope.
6.1.1 Using the OntoUML Dataset
@@
Data: A dataset comprising 10 models developed by the authors of this paper. These are small models (varying from a minimum of 6 classes and 5 relations to a maximum of 10 classes and 8 relations)121212The model files are available here: cmining-approach/tree/main/ontouml/experiment1 that were created specifically for this experiment, taking inspiration from structures that are present in real models, with the goal of reproducing a controlled number of target patterns. The distribution of patterns per model can be observed in Table 1.
Setup: For validation, we used six common OntoUML patterns that served as “litmus test”.131313A litmus test is “a critical indicator of future success or failure” A is a litmus test for B if A can be effectively used to measure some property of B [43]. These patterns were previously manually identified as useful for building OntoUML models by the designers of the language within multiple example models [7, 44]. We represent the selected patterns in Figure 7.
We executed the application seven times (one without adopting parameters to filter specific concepts and the other times with six different configurations, each to find one of the 6 selected target patterns) and we checked whether the proposed solution could discover the pre-identified interesting patterns (R1). Moreover, we tested the role of the customization steps in supporting the discovery process (R2) and the level of comprehensibility of the outcome (R3). First, we conducted a trial adopting no customization facility from the pipeline. As parameters, we selected 3 (the support of the less frequent patterns, i.e., Relator, Subkind, and Phase) as minimum support, and 5 as minimum number of nodes (the size of the smaller pattern, i.e., Characterization),141414Where we have 2 nodes for the classes, 1 node as an association and 2 nodes for the cardinalities. so that all patterns can be discovered. For each other trial, we customized the pipeline to find the target pattern, e.g., by filtering some constructs and some other patterns we did not want to discover. For instance, for the Relator pattern, we selected Kind, Relator, and Role as target constructs and we filtered out information about the generalizations sets (the customization we adopted in terms of minimum support, nodes number, and selected/removed constructs is reported in Table 2.).151515To provide context on the OntoUML constructs present in the patterns, here is a brief explanation: A Kind is a construct commonly used across models to represent rigid concepts that establish an identity principle. A Subkind represents rigid specializations of identity providers, such as Kinds, while the Relator construct is used to represent truth-makers of material relations—entities that must exist for two or more individuals to be connected by material relations. A Role represents anti-rigid specializations of identity providers, like Kinds. The Category construct serves as a rigid mixin, which does not depend on a specific identity principle but is used to aggregate essential properties across individuals with different identity principles. A RoleMixin is the equivalent of Role for types that aggregate instances with different identity principles. The Phase stereotype represents anti-rigid subtypes of identity providers, such as Kinds, that arise due to changes in intrinsic properties (for example, a person’s age). Finally, Characterization is a relation that connects a bearer type with its features. More detailed information on OntoUML constructs can be found in [39] and https://github.com/OntoUML/ontouml-models.
@@
| trial | freq. | nodes | filter | freq. strs. |
@@
| Relator | 3 | 12 | { Select: Kind, Role, Relator } { Remove: Charact } | 42 |
| RoleMixin | 4 | 10 | { Select: Kind, Role, Rmixin } { Remove: Assoc } | 2 |
| Charact. | 8 | 4 | { Select: Kind, Mode } {Remove: Gen } | 1161616Note that the result “1” does not mean that only one occurrence of that pattern was found, but that exactly one structure corresponding to that pattern was found. This structure, in turn, can occur multiple times, for example, in this case, the frequency was “8”, as expected. |
| Category | 6 | 4 | { Select: Kind, Category } { Remove: Assoc } | 32 |
| Subkind | 3 | 4 | { Select: Kind, Subkind } { Remove: Assoc } | 41 |
| Phase | 3 | 4 | { Select: Kind, Phase } { Remove: Assoc } | 89 |
Results: The conducted trial successfully identified all the target patterns, demonstrating the pipeline’s proficiency in addressing R1. The approach identified all the (classes of) patterns that were explicitly identified a priori and correctly counted their occurrences.
@@
6.1.2 Using the ArchiMate Dataset
Similarly to the OntoUML experiment, we used an ad hoc set of ArchiMate models, containing various known patterns as input, and we conducted different trials to verify that the expected patterns can be found in the mining output.
Data: This experiment used a synthetic dataset of 10 manually created ArchiMate models, with sizes ranging from 10 to 15 elements and 8 to 12 relationships. Six EA Smells were selected as target anti-patterns for this analysis, as illustrated in Figure 9. These patterns include Chatty Service (CS), Combinatorial Explosion (CE), Cyclic Dependency (CD), Data Service (DS), Multifaceted Abstraction (MA), and Wrong Cuts (WC). EA Smells [45] serve as qualitative indicators of structural inefficiencies and represent potential issues that affect the non-functional aspects of EA models (e.g., maintenance). Analogous to code smells that signal technical debt in source code, EA Smells assesses an organization holistically, beyond purely technical scopes [45]. The EA Smells used in this experiment were selected from an extensive catalog171717https://swc-public.pages.rwth-aachen.de/smells/ea-smells/ and translated into ArchiMate patterns (cf. [46]) using the provided descriptions. The distribution of the patterns across the dataset is listed in Table 3.
| Model | CS | CE | CD | DS | MA | WC |
| 01 | 1 | 0 | 1 | 0 | 1 | 0 |
| 02 | 0 | 1 | 1 | 0 | 0 | 2 |
| 03 | 0 | 1 | 0 | 1 | 0 | 1 |
| 04 | 1 | 0 | 0 | 1 | 1 | 0 |
| 05 | 0 | 1 | 0 | 0 | 1 | 2 |
| 06 | 0 | 0 | 2 | 0 | 0 | 1 |
| 07 | 0 | 0 | 1 | 0 | 1 | 1 |
| 08 | 1 | 0 | 1 | 1 | 0 | 0 |
| 09 | 0 | 1 | 0 | 0 | 2 | 1 |
| 10 | 0 | 0 | 2 | 0 | 0 | 1 |
| Overall Freq. | 3 | 4 | 8 | 3 | 6 | 9 |
| Model Freq. | 3 | 4 | 6 | 3 | 5 | 7 |
Setup: The experiment was organized into seven trials. The first, referred to as the “neutral” trial, mined patterns without applying any filters to capture all potential patterns. Parameters were set to identify patterns with a frequency of at least 3 (matching the frequency of less frequent patterns, e.g., Chatty Service and Data Service) and a minimum node count of 5 (corresponding to the size of the smallest patterns, e.g., Multifaceted Abstraction and Wrong Cuts).
The subsequent six trials focused on identifying the individual patterns by applying specific filters to narrow the search space. For instance, Cyclic Dependency patterns in the dataset exclusively consist of TechnologyService elements and Triggering relationships. Hence, filters were applied to include only these elements and relationships while excluding others. The mining parameters, such as the minimum frequency and node count, were also adjusted to ensure patterns of interest were captured. The complete configuration of parameters and filters for each trial is summarized in Table 4.
Results: Table 4 summarizes the results of the seven trials. The neutral trial yielded 80 total patterns, including unrelated or redundant outputs, showcasing the need for customized filtering to refine results. In contrast, the filtered trials successfully identified all target patterns, with each trial isolating the intended structure and significantly reducing the number of output patterns (R2). For instance, the Wrong Cuts (WC) pattern, distributed 9 times across 7 models (see Table 3), was correctly identified as a single recurring pattern during the corresponding trial, when applying the specified parameters and filters. Similarly, all other filtered trials correctly identified the intended EA Smell patterns, confirming the approach’s reliability in detecting known structures and accurately counting their occurrences (R1). Note that in the Cyclic Dependency (CD) trial 3 patterns were found, consisting of our target pattern and two duplicates, due to the cycle in the graph structure. An example visualization for a concrete pattern, as returned from the pipeline, is shown in Figure 10.
| trial | freq | nodes | filter | freq. strs. |
| Neutral | 3 | 5 | none | 80 |
| CS | 3 | 9 | { BusinessProcess } { Flow, Serving, Triggering } | 1 |
| CE | 4 | 9 | { TechnologyService, Artifact, ApplicationService } { Realization, Access } | 1 |
| CD | 6 | 8 | { TechnologyService } { Triggering } | 3 |
| DS | 3 | 7 | { ApplicationService, BusinessService, DataObject } { Access, Serving } | 1 |
| MA | 5 | 5 | { BusinessActor, BusinessProcess } { Assignment } | 1 |
| WC | 7 | 5 | { ApplicationComponent, ApplicationFunction } { Realization } | 1 |
6.1.3 Threats to Validity
This experiment focused on a limited set of patterns that, although relatively intricate, represent only a fraction of a broader spectrum of patterns. Moreover, the synthetic datasets used in this study were fairly compact. However, the combination of the patterns we selected from the OntoUML and the ArchiMate scenarios is suitable for assessing whether our approach can discover structures with the required level of expressiveness.
@@ 6.2 Experiment 2: Performance Test
This experiment primarily addresses RQ2. Here, performance is not measured by the number of patterns identified versus those expected, as the chosen mining algorithm has been verified to be 100% accurate [14], making metrics like precision and recall irrelevant. Likewise, the main focus is not the time required to generate output, given that FSM tasks are known to be time-consuming. Instead, the goals are: .i to assess whether the algorithm can produce results on real datasets containing models of varying sizes when run on basic, affordable hardware, and .ii to observe how the execution time scales with an increase in input models.
It is important to note that performance is influenced by the specific implementation choices made for each module in the pipeline, which can vary. For instance, future versions of the application may adopt a different mining algorithm if needed. This highlights the modular nature of our approach, which allows different components to be interchanged to support the same tasks effectively.
6.2.1 Using the OntoUML Dataset
@@
Setup: This validation was executed on a MacBook Pro (Retina, 13-inch, Early 2015) with CPU 2,7 GHz Intel Core i5 and was organized in two main trials: a baseline trial (0), where we selected as input the 47 conceptual models from the OntoUML dataset and trial (1) where we added 47 more models. The selected models have different sizes, in terms of relations and and classes. As an example, the graphs generated from these sources can have a small number of nodes and relations, namely 47 and 35 respectively (see lindeberg2022simple-ontorights.json, in the reference experiments repository), or higher, namely 2.449 and 2.282 (see indeberg2022full-ontorights.json). For each step, to increase the amount of data to be handled, we also tested the approach with six different customizations (we show these values in Table 5) concerning partition size parameters and the minimum frequency of the mining task.
@@
| models | import (s) | nodes | freq. | mining (s) | patterns | clustering (s) | viz. (s) |
| 47 | 0.7327 | 5 | 20 | 1.375 | 1 | 0.011 | 0.106 |
| 3 | 20 | 1.591 | 15 | 0.043 | 1.648 | ||
| 5 | 15 | 3.792 | 22 | 0.073 | 2.257 | ||
| 3 | 15 | 3.950 | 65 | 0.422 | 6.669 | ||
| 5 | 10 | 47.579 | 246 | 0.982 | 16.387 | ||
| 3 | 10 | 48.448 | 372 | 1.360 | 20.387 | ||
| 94 | 1.238 | 5 | 20 | 130.458 | 37 | 0.103 | 3.781 |
| 3 | 20 | 184.239 | 72 | 0.503 | 7.360 | ||
| 5 | 15 | 224.782 | 141 | 0.734 | 9.410 | ||
| 3 | 15 | 229.138 | 193 | 0.812 | 11.500 | ||
| 5 | 10 | 324.781 | 586 | 1.990 | 25.675 | ||
| 3 | 10 | 359.318 | 662 | 3.156 | 31.057 |
6.2.2 Using the ArchiMate Dataset
Data: For this experiment, we used subsets of 50 and 100 medium-sized ArchiMate models, extracted from the EAModelSet dataset [16]. The EAModelSet is a FAIR dataset comprising over 900 ArchiMate models of varying sizes, collected from GitHub, GenMyModel, and the EA community. From the dataset, we first filtered for English-language models and then selected medium-sized models with 30 to 80 relationships each. The 50-model subset contained 2.862 elements and 3.570 relationships, while the 100-model subset comprised 5.890 elements and 6.920 relationships.
Setup: The experiment was run on the machine adopted also in the experiment with OntoUML models (see Section 6.2.1) Two primary trials were conducted: one with 50 models and the other with 100 models. For each trial, six different parameter combinations were tested, progressively increasing the amount of data to be processed. No filters were applied, meaning all elements and relationships in the models were analyzed. A timeout of 600 seconds (10 minutes) was imposed for the mining step to ensure execution feasibility. The parameter combinations and results are presented in Table 6.
Results: Table 6 summarizes the results, detailing the time taken for each step of the pipeline: graph partitioning (import), mining, clustering, and visualization. For the visualization step, only the time required to generate PlantUML.txt files was measured. As expected, the mining step accounted for most of the execution time. The results show that processing time increased under the following conditions:
.i
when the number of total elements and relationships is increased (e.g., moving from 50 to 100 models),
.ii
when the minimum node parameter decreased, allowing smaller patterns to be considered,
.iii
when the frequency threshold was reduced, resulting in a larger number of patterns being mined.
For the 50-model subset, all parameter combinations were completed successfully within the timeout limit. However, for the 100-model subset, the mining step exceeded the 600-second timeout in both trials with the lowest frequency (10), indicating that processing models of this size with such settings would require more time and more robust hardware. These results reaffirm the high computational cost of the mining step, particularly when handling larger datasets with more lenient parameters. The results also show the need to balance parameter settings with hardware capabilities to ensure efficient execution.
| models | import (s) | nodes | freq. | mining (s) | patterns | clustering (s) | viz. (s) |
| 50 | 0.066 | 5 | 20 | 0.178 | 1 | 0.545 | 0.153 |
| 3 | 20 | 0.182 | 11 | 0.562 | 0.137 | ||
| 5 | 15 | 1.257 | 42 | 0.571 | 0.124 | ||
| 3 | 15 | 1.283 | 79 | 0.582 | 0.147 | ||
| 5 | 10 | 66.810 | 502 | 0.593 | 0.159 | ||
| 3 | 10 | 87.120 | 604 | 0.624 | 0.187 | ||
| 100 | 0.154 | 5 | 20 | 125.447 | 8 | 0.657 | 0.201 |
| 3 | 20 | 131.459 | 38 | 0.683 | 0.207 | ||
| 5 | 15 | 455.807 | 68 | 1.032 | 0.314 | ||
| 3 | 15 | 507.853 | 93 | 1.154 | 0.462 | ||
| 5 | 10 | 600+ | - | - | - | ||
| 3 | 10 | 600+ | - | - | - |
6.2.3 Threats to Validity
The primary challenge to the validity of this experiment stems from the absence of a comparative analysis involving multiple devices. The ultimate performance might be influenced by additional variables not present in the current setting. Nonetheless, the configuration we employed can be viewed as a stress test, given its resemblance to the common features found in widely utilized laptops and the absence of a high-performance CPU.